Night Plan UT 29-Dec-2020



Split Night: 1st half LBTI, 2nd half D/E (this plan)

Night Center 00:21

Handover 00:16

Morning Astronomical Twilight: 05:52

Night called for Snow - Closed Dome ESM testing

Night Plan

  • Closed Dome Activities
    • ESM with ARGOS Cal Unit
      • Test ESM TTM table
      • Take test exposures with LUCI

ESM Closed Dome testing

UT 05:07 - beginning ESM testing

L1 0004-0005 2.51 sec darks (for bad pixel mask)

1. Find a rotator position for Bin 2 ESM that does not rail the WFS camera lens.

Rotator = 161 degrees works

2. Take open loop ("seeing limited") data with LUCI-1 and then take closed loop with 1" median seeing to compare improvements - 400Hz

- step through each mag bin

3. Take open loop ("seeing limited") data with LUCI-1 and then take closed loop with 1" median seeing to compare improvements - 100 Hz

- step through each mag bin

4. Repeat for different seeing (seeing is placed on the secondary to simulate 0.5 micron as measured by the DIMM)

Plan for sampling:

0.8" and 1.5" seeing for each, open loop and closed for each and no seeing

Bin 1 R=10.0

Bin 2 R=12.5

Bin 2 R=14.0

Bin 4 R=15.0

Bin 4 R=16.5

UT 05:57 - test exposure R=10 14.7k; peak H-band, there is an internal reflection but it is far enough from the source (2.51sec)

UT 05:58 - test exposure R=16.5-17.0; 2.51 sec gets 32 counts in the peak, trying 30sec --> 540 counts peak, bkgd 130 counts; see an odd "tail" to the left of the star

-turning off source to check if tail vanishes (or if its persistence)

L1 0006-0007 - 30 second H-band - ARGOS cal unit light off

L1 0008-0009 - 2.51 second H-band - ARGOS cal unit light off

200 Hz R=17

L1 0010-0011 - 30sec R=17.0, 200 Hz, seeing=0.75" FWHM= 0.46", 0.48"

L1 0012-0013 -30sec R=17.0, 200 Hz, loop open, no seeing applied, FWHM = 1.37" (out of focus), 1.36" (out of focus) - they look the same - DO NOT USE -bad shape left on AdSec for 0012-0013?

L1 0014-0015 - 30sec R=17.0, 200Hz, seeing=1" FWHM=0.51", 0.56"

L1 0016-0017 - reoptimized gain, 30sec R=17.0, 200Hz, Seeing=1" FWHM=0.51",0.51"

L1 0018-0019 - 30sec R=17.0, 200Hz, no seeing FWHM=0.34", 0.35"

100 Hz R=17

L1 0020-0021 - 30sec R=17.0, 100Hz, no seeing FWHM=0.33",0.32"

L1 0022-0023 - 30sec R=17.0, 100Hz, seeing=0.75" FWHM=0.40",0.41"

L1 0024-0025 - 30sec R=17.0, 100Hz, seeing=1" FWHM=0.48", 0.50"

UT 07:34 - moving to bright end - put blind filter in LUC to reduce persistence - slow bandwidth impacting AO operations

500 Hz R=10

L1 0026-0028 - 2.51sec R=10.0, 500Hz seeing=1" FWHM=0.47",0.48",0.46"

L1 0029-0031 - 2.51sec R=10.0, 500Hz, seeing=0.75" FWHM=0.40",0.41"0.41"

L1 0032-0034 - 2.51 sec R=10.0, 500Hz no seeing, FWHM=0.33",0.33",0.33"

100 Hz R=10

L1 0035-0037 - 2.51 sec R=10.0 100Hz, no seeing, FWHM=0.32",0.33",0.36"

L1 0038-0040 - 2.51 sec R=10.0, 100Hz, seeing=0.75" FWHM=0.43",0.40",0.41"

L1 0041-0043 - 2.51 sec R=10.0, 100Hz, seeing=1" FWHM=0.51",0.45",0.51"

400 Hz R=12.5

L1 0044-0046 - 15sec R=12.5, 400Hz, em_gain=600, seeing=1" FWHM=0.47",0.47",0.48"

L1 0047-0049 - 15sec R=12.5, 400Hz, em_gain=600 seeing=0.75" FWHM=0.42",0.41",0.41"

L1 0050-0052 - 15sec R=12.5, 400Hz, em_gain=600, no seeing FWHM=0.31",0.28",0.29"

100 Hz R=12.5

L1 0053-0055 - 15sec R=12.5, 100Hz, no seeing 0 gains FWHM=0.34,0.32",0.33" - not sure this is useful, mainly shows if applying gains is working- and it does

L1 0056-0058 - 15sec R=12.5, 100Hz optimize gains em_gain=600, no seeing FWHM=0.33",0.33",0.32

L1 0059-0061 - 15sec R=12.5, 100Hz, em_gain changed from 600 to 100, puts fewer photons on subapertures 0.75"seeing - FWHM=0.40",0.41",0.40"

L1 0062-0064 - 15sec R=12.5, 100Hz, em_gain=100, 1" seeing, FWHM=0.45",0.47",0.45"

L1 0065-0067 - 15sec R=12.5, 100Hz em_gain=100, 0.75"seeing FWHM=0.40",0.39",0.40"

L1 0068-0070 - 15sec R=12.5 100Hz, em_gain=100, no seeing FWHM= 1st image is a triangle, all images triangles - DO NOT USE

L1 0071-0073 - 15sec R=12.5 100Hz em_gain=100, no seeing FWHM=0.30",0.31",0.33"

400 Hz R=14.9

L1 0074-0076 - 30sec R=14.9, 400Hz no seeing FWHM=0.30",0.29",0.29"

L1 0077-0079 -30sec R=14.9, 400Hz 0.75" seeing FWHM=0.40",0.40",0.40"

L1 0080-0082 -30sec R=14.9, 400Hz, 1" seeing FWHM=0.46",0.47",0.48"

100 Hz R=14.9

L1 0083-0085 - 30sec R=14.9, 100Hz em_gain=300, 1" seeing FWHM=0.75",0.76"- DO NOT USE - needed to optimize gain

L1 0086-0088 - 30sec R=14.9 100Hz em_gain=300, 1" seeing FWHM=0.53",0.50",0.49"

L1 0089-0091 - 30sec R=14.9 100Hz em_gain=300, 0.75" seeing FWHM = 0.44",0.44",0.44"

L1 0092-0094 - 30sec R=14.9 100Hz em_gain-300 no seeing FWHM = 0.35",0.33",0.33"

400 Hz R=13.9

L1 0095-0097 - 20sec R=13.9 400Hz no seeing FWHM=0.24",0.24",0.24" - DO NOT USE WRONG RECONSTRUCTOR

-why would the wrong reconstructor give a better PSF??

L1 0098-0100 - 20sec R=13.9 400Hz no seeing FWHM=0.32",0.32",0.32"

L1 0101-0103 - 20sec R=13.9 400Hz 0.75" seeing FWHM=0.42",0.42",0.41"

L1 0104-0106 - 20sec R=13.9 400Hz 1" seeing FWHM = 0.48",0.49",0.50"

100 Hz R=13.9

L1 0107-0109 - 20sec R=13.9 100Hz 1"seeing FWHM=0.50", 0.49",0.49"

L1 0110-0112 - 20sec R=13.9 100Hz 0.75" seeing FWHM=0.42",0.42",0.42"

L1 0113-0115 - 20sec R=13.9 100Hz no seeing FWHM=0.31",0.31",0.31"

L1 0116-0117 20 sec background (ARGOS lamp off)

L1 0118-0119 15 sec background (ARGOS lamp off)

UT 13:02 - complete LUCI in safe config

-- BarryRothberg - 28 Dec 2020
Topic revision: r3 - 29 Dec 2020, BarryRothberg
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback